Circular No. 6877

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ar IN NGC 2916 Y. L. Qiu, Q.Y. Qaio, W. D. Li, and J. Y. Hu, Beijing Astronomical

Observatory (BAO) report the discovery of an apparent supernova on unfiltered CCD frames taken in the course of the BAO Supernova Survey on Apr. 14 and 15 UT, when the estimated magnitudes of the new star were 18.4 and 18.2, respectively. SN 1998ar is located at α = 9^{h}34^{m}58^{s}.86, δ = +21^{o}42′57′′.5 (equinox 2000.0), which is 16′′.9 east and 38′′.5 north of the center of NGC 2916. CCD frames taken on Mar. 14 (limiting mag ∼ 18.6) show no star at the position of SN 1998ar.

SAX J1808.4–3658 = XTE J1808–369 D. Chakrabarty and E. H. Morgan, Massachusetts Institute of Technol-

ogy, report for the Rossi XTE team: “The pulsations from XTE J1808–369 (= SAX J1808.4–3658) recently discovered in RXTE observations (IAUC 6876) show clear evidence for a binary orbit. Using RXTE/PCA data from Apr. 11.81–11.82, 11.86–11.89, and 13.07–13.09 and assuming a constant pulsar-spin frequency for all these data, we find the following provisional orbital parameters for the neutron star: an orbital period of 7251(1) s, a projected semimajor axis of 0.0632(3) lt-s, and an orbital epoch (time of 90^{o} mean longitude) of JD 2450915.3995(1). The corresponding barycentric spin frequency is 400.9753(1) Hz. The orbital parameters give a pulsar mass function of 3.9 × 10^{−5} M_{⊙}, one of the smallest mass functions for any x-ray binary. Since x-ray heating of the mass donor and/or the accretion disk in such a low-mass x-ray binary is likely to be very strong, observations at other wavelengths are urgently required to identify the counterpart while the x-ray source is still active.”

SUPERNOVA 1998an IN UGC 3683 D. Hanzl, N. Copernicus Observatory and Planetarium, Brno, reports

the following position for SN 1998an (cf. IAUC 6871) obtained with a 0.4-m reflector (+ CCD + V filter) on Apr. 11.8 UT: α = 7^{h}08^{m}17^{s}.13 ± 0^{s}.02, δ = +46^{o}06′54′′.3 ± 0′′.8 (equinox 2000.0). Hanzl also provides the following CCD V magnitudes (based on GSC comparison stars): Apr. 8.792, 15.6; 11.819, 15.6; 13.815, 15.7; 14.856, 15.6.

1998 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green